# Fortify the Shortest Stave in Attention: Enhancing Context Awareness of Large Language Models for Effective Tool-Use

### v2

@@ Abstract
In this paper, we demonstrate that an inherent waveform pattern in the attention allocation of large language models (LLMs) significantly affects their performance in tasks demanding a high degree of context awareness, such as utilizing LLMs for tool-use. Specifically, the crucial information in the context will be potentially overlooked by model when it is positioned in the trough zone of the attention waveform, leading to decreased performance. To address this issue, we propose a novel inference method named Attention Buckets. It allows LLMs to process their input through multiple parallel processes. Each process utilizes a distinct base angle for the rotary position embedding, thereby creating a unique attention waveform. By compensating an attention trough of a particular process with an attention peak of another process, our approach enhances LLM’s awareness to various contextual positions, thus mitigating the risk of overlooking crucial information. In the largest tool-use benchmark, our method elevates a 7B model to achieve state-of-the-art performance, comparable to that of GPT-4. On other benchmarks and some RAG tasks, which also demand a thorough understanding of contextual content, our Attention Buckets also exhibited notable enhancements in performance. [^1] We release our code at https://github.com/AlibabaResearch/DAMO-ConvAI/attention-buckets.
@@ 1 Introduction
Figure 1: (a) Task illustration: Presented with multiple key-value pairs and a target key (highlighted in bold), the model is required to accurately retrieve and generate the value associated with this key from an extensive context. (b) We illustrate the position-related fluctuation in accuracy of Llama-2-7B on this in-context retrieval task. (c) The upper bound of the attention score exhibits fluctuations, which we term the “attention waveform”. Our study reveals a connection between the position-related fluctuations in LLMs’ performance and this attention waveform.
Despite the achievements made by current LLM-based tool-use frameworks, in our practical experience, we observed that LLMs exhibit varying levels of awareness concerning different positions within the context. For instance, LLMs may overlook certain tools within the context, resulting in a failed call; however, by altering the position of these tools, the task can be successfully executed. Such variations significantly affect the performance of LLMs in tool-use. This observation is consistent with the findings from a previous study [9] that investigated a simple in-context retrieval task. When LLMs are presented with multiple key-value pairs and instructed to retrieve the value associated with a specific key, the index of the queried target key results in significant fluctuations in accuracy. Figure 1(a) provides a visual representation of the instructions for this task. Figure 1(b) shows this fluctuation we replicated using the Llama-2-7B [10]. In our study, we go beyond the superficial fluctuations previously observed and identify that these position-related performance differences are closely associated with the model’s fluctuating attention allocation. Specifically, we noticed that the upper bound of the attention score exhibits a waveform pattern, as depicted in Figure 1(c). We demonstrate that if the position of the crucial information coincides with a trough in the attention waveform, the model may overlook it, leading to decreased accuracy.
@@
(1) For LLMs with RoPE, we propose and verify an explanation for the variation in their awareness of different positions within the context. We establish a relationship between this variation and the attention waveform, i.e., the fluctuation in the upper bound of attention score.
@@ 2.1 Preliminaries
When the relative distance $n-m$ increases, the upper bound of the attention score before softmax demonstrates a long-term decay, i.e., the value generally decreases as the relative distance grows. This trend is accompanied by a waveform, as depicted in Figure 1(c). The derivation of the upper bound is presented in the appendix A.
@@ 2.2 Hypothesis Verification
We varied the RoPE base within the model from $10,000$ to $30,000$ in increments of $5,000$. For each base value, we calculate the corresponding upper bound of attention score and identify the positions of the peaks and troughs (see appendix B for details). Each test sample undergoes two evaluation rounds: In the first round, we position the target key-value pair at the attention peak nearest to the exact middle of the context. In the second round, we move the target pair to the nearest attention trough. By comparing accuracy differences between the two rounds, we aimed to answer how much attention waveform patterns impact the model’s context awareness. The experiment was conducted with varying context lengths by setting different $K$ (40 and 50, respectively). We provide more experimental details in appendix C.
@@ 3 Enhancing Context Awareness via Interleaving Attention Waveform
Based on the insights presented above, we introduce a novel approach to sidestepping the above challenge, with the goal of improving the LLM performance in tool-use by enhancing its context awareness. Our method focuses on the inference stage of LLMs and does not require training. We first provide preliminary definitions, followed by a detailed introduction to our approach.
@@ 3.2 Method
[…]
@@ 4 Experiments
[…]
@@ 4.1 Experiment Setups
[…]
@@
[…]
@@ 4.3 Discussion on Efficacy
Original +$\text{AB}_{once}$ +USC +AB
pass win pass win pass win pass win
I1-Inst. 64.0 62.3 52.0 37.0 66.0 63.0 68.5 65.0
I1-Cat. 64.0 59.0 40.5 31.0 65.5 61.5 70.0 65.5
I1-Tool. 60.5 55.5 47.0 34.5 61.0 58.5 65.0 67.0
I2-Inst. 81.5 68.5 70.5 65.0 80.0 73.0 84.0 78.0
I2-Cat. 68.5 60.8 65.0 58.0 70.0 61.5 71.0 64.5
I3-Inst. 65.0 73.0 61.0 52.0 66.0 78.0 69.0 89.0
Avg 67.3 63.1 56.0 45.3 68.1 65.9 71.3 71.5
Table 3: Comparison among Attention Buckets, $\textit{Attention Buckets}_{once}$, and USC, based on the ToolLlama-DFSDT-Retriever configuration. We use the abbreviation AB for Attention Buckets.
$\bullet$ Universal Self-Consistency. USC [27] is a universal self-consistency [28] algorithm that supports free-format outputs. The LLM first generate $N$ responses. Subsequently, the LLM is tasked with selecting the response that exhibits the highest degree of consistency, employing a specific prompt. The memory cost of this method is roughly equivalent to that of our Attention Buckets.
@@
USC incurs a similar inference cost to our method. However, it only demonstrates a mere 0.8% point increase in pass rate and a 2.8% point increase in win rate on average, compared to the original ToolLlama. This limited enhancement can be attributed to the USC method’s failure to effectively address the problem of trough position oversight inherent in an inference with a single RoPE base. Despite multiple attempts, this oversight persists. This comparison clearly illustrates the effectiveness of the Attention Buckets  as it outperforms the method with a similar level of overhead and without additional training.
@@ 5 Exploring Applications for Retrieval-Augmented Generation
We use Llama-2-7B [10] as our backbone model and compare it with FiD-XL [34], the exclusive ODQA model trained on multiple ODQA benchmarks, including NQ and WebQ. The results are shown in Table 4. When augmented with Attention Buckets, Llama-2-7B demonstrated superior performance over the dedicated QA model FiD-XL.
Method NQ WebQA
FiD-XL (3B) 50.1 50.8
Llama-2 (7B) 48.5 51.7
+ Attention Buckets 50.3 (1.8$\uparrow$) 53.1(1.4$\uparrow$)
Table 4: Accuracy on NQ and WebQ (10 documents).
@@ 8 Conclusion
In this paper, we delved into the waveform patterns observed in attention scores and found that fluctuations in the upper bound of attention score could potentially affect the model’s context awareness, particularly in relation to the position of crucial information within the context. We propose Attention Buckets, an inference augmentation method designed to enhance the model’s context awareness. This method combines various attention patterns, which are controlled by different RoPE bases. Our approach has achieved state-of-the-art performance on the current largest tool-use benchmark while showing the applicability to a wider range of RAG tasks.
@@ Acknowledgement
Appendix A The Upper Bound of Attention Score Before Softmax
@@
When $q_{m}$ and $k_{n}$ are identical, the inner-product reaches its maximum value. For computational simplicity, we assume them as all-one vectors to derive the upper bound (UB) of the inner-product (attention score before softmax):
$UB=\sum^{d/2-1}_{j=0}2\cos\left((m-n)\theta_{j}\right)\geq q_{m}\cdot k_{n}.$ (7)
Figure 1(c) illustrates the visualization of $UB$ with base = 10,000. Figure 3(b) results from varying base with values from our searched set $\mathcal{B}_{c}$. These figures demonstrate the horizontal axis as the relative position between $k_{n}$ and $q_{m}$.
@@ Appendix B Locating Peaks and Troughs in an Attention Waveform
# Calculate the upper bound of attention score before softmax.
@@
    # return: the upper bound of attention score between qm and kn.

### v4

@@ Abstract
In this paper, we demonstrate that an inherent waveform pattern in the attention allocation of large language models (LLMs) significantly affects their performance in tasks demanding a high degree of context awareness, such as utilizing LLMs for tool-use. Specifically, the crucial information in the context will be potentially overlooked by model when it is positioned in the trough zone of the attention waveform, leading to decreased performance. To address this issue, we propose a novel inference method named Attention Buckets. It allows LLMs to process their input through multiple parallel processes. Each process utilizes a distinct base angle for the rotary position embedding, thereby creating a unique attention waveform. By compensating an attention trough of a particular process with an attention peak of another process, our approach enhances LLM’s awareness to various contextual positions, thus mitigating the risk of overlooking crucial information. In the largest tool-use benchmark, our method elevates a 7B model to achieve state-of-the-art performance, comparable to that of GPT-4. On other benchmarks and some RAG tasks, which also demand a thorough understanding of contextual content, our Attention Buckets also exhibited notable enhancements in performance. [^1] We release our code at https://github.com/Fiorina1212/Attention-buckets.
@@ 1 Introduction
Figure 1: (a) Task illustration: Presented with multiple key-value pairs and a target key (highlighted in bold), the model is required to accurately retrieve and generate the value associated with this key from an extensive context. (b) We illustrate the position-related fluctuation in accuracy of Llama-2-7B on this in-context retrieval task. (c) The pattern of the attention score exhibits fluctuations, which we term the “attention waveform”. Our study reveals a connection between the position-related fluctuations in LLMs’ performance and this attention waveform.
Despite the achievements made by current LLM-based tool-use frameworks, in our practical experience, we observed that LLMs exhibit varying levels of awareness concerning different positions within the context. For instance, LLMs may overlook certain tools within the context, resulting in a failed call; however, by altering the position of these tools, the task can be successfully executed. Such variations significantly affect the performance of LLMs in tool-use. This observation is consistent with the findings from a previous study [9] that investigated a simple in-context retrieval task. When LLMs are presented with multiple key-value pairs and instructed to retrieve the value associated with a specific key, the index of the queried target key results in significant fluctuations in accuracy. Figure 1(a) provides a visual representation of the instructions for this task. Figure 1(b) shows this fluctuation we replicated using the Llama-2-7B [10]. In our study, we go beyond the superficial fluctuations previously observed and identify that these position-related performance differences are closely associated with the model’s fluctuating attention allocation. Specifically, we observed a waveform pattern in the attention “intensity” (referred to as the attention waveform in this paper) when LLMs retrieve the same token from the context, as illustrated in Figure 1(c). We demonstrate that if the position of the crucial information coincides with a trough in the attention waveform, the model may overlook it, leading to decreased accuracy.
@@
(1) For LLMs with RoPE, we propose and verify an explanation for the variation in their awareness of different positions within the context. We establish a relationship between this variation and the attention waveform.
@@ 2.1 Preliminaries
When the relative distance $n-m$ increases, the waveform of the attention score before softmax demonstrates a long-term decay, i.e., the value generally decreases as the relative distance grows. This trend is accompanied by a waveform, as depicted in Figure 1(c). The derivation of this waveform is presented in Appendix A.
@@ 2.2 Hypothesis Verification
We varied the RoPE base within the model from $10,000$ to $30,000$ in increments of $5,000$. For each base value, we calculate the corresponding waveform of attention score and identify the positions of the peaks and troughs (see Appendix B for details). Each test sample undergoes two evaluation rounds: In the first round, we position the target key-value pair at the attention peak nearest to the exact middle of the context. In the second round, we move the target pair to the nearest attention trough. By comparing accuracy differences between the two rounds, we aimed to answer how much attention waveform patterns impact the model’s context awareness. The experiment was conducted with varying context lengths by setting different $K$ (40 and 50, respectively). We provide more experimental details in appendix C.
@@ 3 Enhancing Context Awareness via Interleaving Attention Waveform
Based on the insight presented above, we introduce a novel approach to sidestepping the above challenge, with the goal of improving the LLM performance in tool-use by enhancing its context awareness. Our method focuses on the inference stage of LLMs and does not require training. We first provide preliminary definitions, followed by a detailed introduction to our approach.
@@ 3.2 Method
Given an input context $\mathcal{C}$, our approach involves duplicating this context into $N$ copies, forming a batch that allows for parallel processing by the LLM. In each parallel, each of these $N$ copies is individually processed with a distinct RoPE base $B_{j}$ from a base set $\mathcal{B}_{c}$, resulting in $N$ corresponding predicted distribution $p$ over the vocabulary $V$. Our selection of $\mathcal{B}_{c}$ guarantees that an attention trough in one parallel is compensated by a peak in another, effectively reducing the possibility of the LLM missing essential information residing within an attention trough. We will delve into the details of determining $\mathcal{B}_{c}$ in §$\S$ 3.3.
@@ 4 Experiments
Figure 3: (a) RoPE Base Searching: we measure the distance from the candidate bases’ attention peaks (or troughs) to the attention troughs (or peaks) corresponding to items in $\mathcal{B}_{c}$. For demonstration clarity, we illustrate only a partial context that corresponds to one waveform period. (b) Attention waveform corresponding to $\mathcal{B}_{c}$ searched by hyperparameters detailed in §$\S$3.3.
@@ 4.1 Experiment Setups
The Toolbench evaluation consists of three distinct levels and three specific scenarios, each offering its own set of challenges. The three evaluation levels include Inst.: testing the model’s response to new instructions for tools already covered in the training data; Tool.: measuring performance with unfamiliar tools within the same tool categories as those in the training dataset; and Cat.: examining the model’s ability to handle tools from completely new categories not represented in the training data. The scenarios are I1: single-tool instructions, I2: multi-tool instructions within the same category, and I3: multi-tool instructions spanning across different collections. Due to specific details, there are only six combinations of levels and scenarios: I1-Inst., I1-Tool., I1-Cat., I2-Inst., I2-Cat., and I3-Inst. Each combination comprises 200 test queries, with the exception of I3-Institution, which includes 100 queries. For a more detailed introduction, readers are recommended to  [4, §$\S$3.2].
@@
Based on the training dataset in ToolBench, [4] fine-tuned a model named ToolLlama, building upon Llama-2-7B [10]. The authors compare it with advanced close-source LLMs, including ChatGPT [2], Claude-2, Text-Davinci-003, and GPT-4 [3]. We implement Attention Buckets to enhance the performance of ToolLlama. The $\mathcal{B}_{c}$ employed for this benchmark is the same as detailed in §$\S$3.3 and depicted in Figure 3(b). Following [4], we adopt multiple reasoning methods for each model, including ReACT [25], DFSDT [4, 26], and an API retriever [4] augmentation for reducing noise in tool selection (DFSDT-Retriever). We adopt the greedy decoding strategy.
@@ 4.3 Discussion on Efficacy
$\bullet$ Attention Sorting. In ASort [27], tokens in distant contexts that receive high attention are considered important. The authors first segmented the context and calculated the average attention of each segment. By rearranging segments in context based on sorted attention scores(with the highest attention segment placed last), they generate the answer using the newly sorted context. This approach does not require extra memory; however, it necessitates multiple iterations to gain the attention scores and lacks parallelizability.
$\bullet$ Universal Self-Consistency. USC [28] is a universal self-consistency [29] algorithm that supports free-format outputs. The LLM first generates $N$ responses. Subsequently, the LLM is tasked with selecting the response that exhibits the highest degree of consistency, employing a specific prompt. The memory cost of this method is roughly equivalent to that of our Attention Buckets.
Original +$\text{AB}_{once}$ +USC +ASort +AB
pass win pass win pass win pass win pass win
I1-Inst. 64.0 62.3 52.0 37.0 66.0 63.0 66.0 63.0 68.5 65.0
I1-Cat. 64.0 59.0 40.5 31.0 65.5 61.5 67.0 62.0 70.0 65.5
I1-Tool. 60.5 55.5 47.0 34.5 61.0 58.5 59.5 58.5 65.0 67.0
I2-Inst. 81.5 68.5 70.5 65.0 80.0 73.0 80.0 68.5 84.0 78.0
I2-Cat. 68.5 60.8 65.0 58.0 70.0 61.5 68.5 60.3 71.0 64.5
I3-Inst. 65.0 73.0 61.0 52.0 66.0 78.0 66.0 75.0 69.0 89.0
Avg 67.3 63.1 56.0 45.3 68.1 65.9 67.8 64.6 71.3 71.5
Table 3: Comparison among Attention Buckets(AB), $\textit{Attention Buckets}_{once}$ ($\text{AB}_{once}$), Universal Self-Consistency (USC) and Attention Sorting (ASort), based on the ToolLlama-DFSDT-Retriever configuration.
@@
ASort tries to strengthen the model’s focus on key information through re-sorting, but its weak improvement in task performance, with a 0.5% point increase in pass rate and 1.5% point in win rate on average, reveals that the attention scores may not capture crucial information well.
USC incurs a similar inference cost to our method. However, it only demonstrates a mere 0.8% point increase in pass rate and a 2.8% point increase in win rate on average, compared to the original ToolLlama. This limited enhancement can be attributed to the USC method’s failure to effectively address the problem of trough position oversight inherent in inference with a single RoPE base. Despite multiple attempts, this oversight persists. This comparison clearly illustrates the effectiveness of the Attention Buckets  as it outperforms the method with a similar level of overhead and without additional training.
@@ 5 Exploring Applications for Retrieval-Augmented Generation
Method NQ WebQA
FiD-XL (3B) 50.1 50.8
Llama-2 (7B) 48.5 51.7
+ ASort 48.9 52.1
+ USC 47.6 51.7
+ Attention Buckets 50.3 (1.8$\uparrow$) 53.1(1.4$\uparrow$)
Table 4: Accuracy on NQ and WebQ (10 documents).
@@
We use Llama-2-7B [10] as our backbone model and compare it with FiD-XL [35], the exclusive ODQA model trained on multiple ODQA benchmarks, including NQ and WebQ.
The results are shown in Table 4. Compared with the other two methods, ASort [27] and USC [28], Attention Buckets exhibits a more stable improvement. Additionally, when augmented with Attention Buckets, Llama-2-7B demonstrated superior performance over the dedicated QA model FiD-XL.
@@ 8 Conclusion
In this paper, we delved into the waveform patterns observed in attention scores and found that waveform of the attention score could potentially affect the model’s context awareness, particularly in relation to the position of crucial information within the context. We propose Attention Buckets, an inference augmentation method designed to enhance the model’s context awareness. This method combines various attention patterns, which are controlled by different RoPE bases. Our approach has achieved state-of-the-art performance on the current largest tool-use benchmark while showing the applicability to a wider range of RAG tasks.
@@ Acknowledgement
Appendix A The Waveform of Attention Score Before Softmax
@@
When $q_{m}$ and $k_{n}$ are identical, the inner-product reaches its maximum value. For computational simplicity, we assume them as all-one vectors to derive the waveform ($\mathcal{W}$) of the inner-product (attention score before softmax):
$\mathcal{W}=\sum^{d/2-1}_{j=0}2\cos\left((m-n)\theta_{j}\right)\geq q_{m}\cdot k_{n}.$ (7)
Figure 1(c) illustrates the visualization of $\mathcal{W}$ with base = 10,000. Figure 3(b) results from varying base with values from our searched set $\mathcal{B}_{c}$. These figures demonstrate the horizontal axis as the relative position between $k_{n}$ and $q_{m}$.
@@ Appendix B Locating Peaks and Troughs in an Attention Waveform
# Calculate the waveform of attention score before softmax.
@@
# return: the waveform of attention score between qm and kn.